OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (3, –2) and is parallel to the line y = –4x + 1?

OpenStudy (anonymous):

x + 4y = 10 4x + y = 10 x – 4y = –10 4x – y = –10

OpenStudy (anonymous):